# An Interpretable Deep-Learning Framework for Predicting Hospital Readmissions From Electronic Health Records

### v1

@@ 3.2 Readmission Prediction Model
Our model is based on a neural network that receives as input two distinct vectors, the first containing the temporal succession of hospital admissions (blue dashed box in Fig. 2), the other containing the corresponding time gaps at each admission (red dashed box in Fig. 2). The core of our solution is represented by the Embedding layer and the ConvLSTM architecture.
@@
The Embedding layer receives as input a patient’s medical record for a specific admission in the form of temporally divided ICD-9-CM codes, as represented in Fig. 1. At this point, each code is converted into the corresponding unique word embedding vector, allowing us to better represent their semantics. This layer is necessary since computers work with numbers and cannot process strings in their native format. To improve the performance of the system, we opted for a transfer learning approach. Specifically, the pre-trained word embedding was obtained by [21], who used the word2vec model with SkipGram implementation [19] to train on healthcare records of 4 million patients over a time span of 2-4 year. The embedding vectors are 300-dimensional and contain many medical code standards, in our case we use only ICD-9-CM related to diagnoses and procedures. Using dimensionality reduction techniques such as t-SNE [26], it is possible to represent the vectors generated by word embedding in 2 dimensions. This procedure allows to graphically represent the semantically similar vectors as two neighboring points. Fig. 3 shows the result produced, obtained using t-SNE with “cosine similarity” as evaluation metric to estimate the semantic relatedness. In Fig. 3 we show ICD-9-CM codes of both diagnoses and procedures, colored in blue and orange respectively. As we can see from the graph, the interaction space of diagnoses and procedures is very complex, but it is possible to identify clusters defined by a specific medical domain.
@@
After each code has been converted into the corresponding embedding vectors, the ConvLSTM architecture implements the forecasting algorithm. This deep learning framework, introduced in [27], is an extension of LSTM, in which the matrix multiplication is replaced with the convolution operation. Thanks to its structure, this innovative architecture allows to exploit both the spatial and temporal information of the input, characteristics necessary for our system to consider both long-term dependencies between the admissions and abstract general health information (e.g., comorbidity) from sequences of specific procedures and diagnosis. Following the ConvLSTM layer, max-pooling is performed, with the goal of making data depend only on the number of filters applied in the convolution and not on the initial input size. The second input contains a vector of categorical data (red dashed box in Fig. 2), equal in size to the time frames considered. The categories are those related to the various time gaps. Such data, being a simple sequence of mono-dimensional vectors, are processed by an LSTM layer, which permits to give the right weight to each event passed in the sequence. Subsequently, the outputs coming from the two branches of the network are concatenated (green dashed box in Fig. 2) and then a classifier with two Dense layers is used to predict whether a readmission will take place or not.
@@ 4 Experiments and Results
In this section, we first describe the training data used for ConvLSTM1d, then we report and discuss the performances on two prediction tasks: readmission within 30 and 180 days. The main objective is to verify that our model improves performances in predicting unplanned hospital readmissions compared to state-of-the-art models, while at the same time allowing the interpretability of the results. Moreover, we compare ConvLSTMd1 with Deepr [10] (a CNN-based model for readmission prediction) and with two traditional machine-learning models, namely logistic regression and random forest. The latter models are reported since they are widely used in hospital readmission prediction [28]. For logistic regression and random forest, we use a bag-of-words model instead of word embeddings. Bag of words is a simple yet popular approach in NLP problems, where text is represented as a set of words, without considering semantics or word order [29].
@@ 4.1 Training Dataset
The EHRs used to train ConvLSTM1d are taken from the MIMIC III Database, or simply MIMIC (Medical Information Mart for Intensive Care) [17]. MIMIC is a large, freely-available database comprising anonymized healthcare data associated with over 40,000 patients who stayed in intensive care units. The initial pre-processing of the dataset is fundamental to avoid errors in the data that may propagate to the subsequent steps. During pre-processing, we take inspiration from the procedure described in [9], where the authors also use MIMIC III data for a task of readmission prediction.
@@
Fig. 4a and 4b report the data distributions from which we gathered this information. As a result, the two datasets we are going to use have the following characteristics:
@@
Additionally, the two generated datasets are balanced. Otherwise, as evidenced by Fig. 4c and 4d, the negative label would be predominant in both tasks (with a prevalence of 94% and 87% respectively), and accuracy could not be used as a metric. In both datasets, there may be multiple admissions for a single person. The presence of multiple admissions is important to extrapolate as much information as possible from the rows of the datasets with a positive label (minority label).
@@ 4.2 Training Parameters
After splitting the datasets into training, validation and test sets, we train our model using Adam optimizer and binary cross-entropy loss function. Table 1 reports the hyperparameters used for the two tasks of 30-days and 180-days admission. Fig. 5 represents the learning curves for both loss and accuracy on the 30-day unplanned readmission task.
Table 1: Hyperparameters used for ConvLSTM1d.
@@
4.3 Metrics
Before presenting the performance obtained by ConvLSTMd1, we carry out a brief summary of the main metrics adopted. In the following description, TP stands for True Positive, TN for True Negative, FP for False Positive, FN for False Negative.
Accuracy is an intuitive metric that consists in the ratio between the correctly predicted observations and the total number of observations. To obtain a good validity, the dataset must be balanced. $Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$.
Precision is defined as the ratio of correctly predicted positive observations over total predicted positive observations. It is a good metric when the cost of False Positives is high. $Precision=\frac{TP}{TP+FP}$.
Recall is the ratio of observations correctly classified as positive over the total number of observations that are actually positive. It is a good metric when the cost of False Negatives is high. $Recall=\frac{TP}{TP+FN}$.
F1 Score is the harmonic mean between precision and recall, allowing to take into account both False Positives and False Negatives. $F_{1}=2\cdot\frac{\mathrm{p}\cdot\mathrm{r}}{\mathrm{p}+\mathrm{r}}$, where $p$ stands for Precision and $r$ for Recall.
@@ 4.4 Results and Discussion
Table 2 illustrates the performance results of ConvLSTMd1 for both 30-days and 180-days unplanned readmission prediction. Our model is trained to achieve the highest accuracy, which is used as the main evaluation metric. To give a better overview, we compare the results of ConvLSTMd1 with two traditional machine-learning approaches for data prediction (logistic regression and random forest) and with Deepr [10], the most similar and advanced solution in literature for predicting hospital readmissions from EHRs. All results are obtained from carrying out 10 separate training sessions for each task and considering the average score between sessions. This expedient is necessary since deep-learning models tend to obtain slightly different performance scores even for the same task.
@@
From Table 2 it is evident that ConvLSTM1d is the best model in terms of accuracy, substantially outperforming all the other models. In particular, logistic regression and random forest obtain good results in either precision or recall for certain specific cases, but with overall accuracy and F1 scores significantly lower compared to our model. Deepr considerably improves on both accuracy and F1 scores with respect to traditional machine-learning methods, but it prevails over ConvLSTM1d only in recall for unplanned readmission within 30 days. However, it should be noted that precision is much lower. Consequently, this is reflected in the F1 score, which favors ConvLSTM1d in both tasks. On the other hand, ConvLSTM1d obtains higher results in both accuracy and F1 scores for both labels and both tasks.
5 Interpretability
Deep learning techniques have succeeded, in recent years, to produce state-of-art results in many areas. One of the most critical aspects of these techniques is, however, in the interpretability of the results. Such models are often seen as “black boxes” for which there is no internal transparency. The reason lies in the structure of the networks, which are composed of various layers that apply non-linear transformations.
The interpretability of the results of predictive tasks is of paramount importance in critical domains such as medicine, where a single choice can compromise a person’s life. For this reason, we carried out an analysis of currently available interpretability frameworks. A first attempt at making deep learning models for readmission prediction more interpretable was introduced by the author of [10], where a method based on convolution filters inspection is presented. Convolution filters act as n-grams detectors, which in the case of our model can be seen as “clinical motifs”. As we presented in the introductory part of the paper, detecting clinical motifs is crucial for accurately predicting unplanned readmission. Unfortunately, as subsequently demonstrated by [24], such filters are often difficult to interpret, because a single filter may not only be able to identify multiple n-grams, but it may also cancel the contribution of other n-grams.
In this context, a popular framework is LIME, a technique frequently used to make machine-learning and deep-learning models interpretable. Specifically, we tested LIME to obtain the contribution that each code contained in the EHR has on the overall prediction. The results are obtained by performing a perturbation of the initial input, based on the concept of “local fidelity” [23]. This provides a straightforward solution to understand the determinants of a certain prediction. Unfortunately, this second method also do not yield satisfactory interpretability results. In fact, the input of ConvLSTM1d model has a specific structure—two vectors, one for the ICD-9-CM codes and the other for the time gaps, both structured temporally—that does not allow for easy perturbation.
For these reasons, we decided to develop an ad-hoc model-dependent interpretability procedure capable of delivering satisfactory results when EHRs comprise also temporal information in their structure. The aim of this method is to represent the importance that each code has on the overall results, thus helping the medical professional understand which are the factors that influence the final prediction the most. As presented in [25], such an approach is possible since the architecture of our prediction model is rather simple (composed of just few layers), thus greatly improving its interpretability. The goal is to have an easy-to-understand output structured in the following way:
  • Patient’s EHR, consisting of the various ICD-9-CM codes.
  • The prediction results, with probability of each label.
Figure 6: Implementation of the model-dependant interpretability procedure – first step.
Figure 7: Implementation of the model-dependant interpretability procedure – second step.
We now present the procedure step by step. First, if we consider the output of a generic layer of our neural network, the correspondence with the code that originated this output is lost due to the max-pooling layer. In fact, this layer selects the largest response for each filter. To obtain an interpretable result, the first step to perform is to keep track, for each filter, of the input code that originates the largest response. In doing so, at the end of this step, we obtain a structure that contains, for each input code, the correspondence between a filter and an ICD-9-CM code, here named FILTER-CODE. It may happen that a code is not the maximum response of any filter, therefore not all the codes will be present in the FILTER-CODE structure. At this point, we have the correspondence between the output of the max-pooling layer and the input codes.
The second step is to modify the intermediate results coming out of the max-pooling layer. The results of the max-pooling layer are necessarily $\geq 0$, since this layer is preceded by a ReLu activation function. Consequently, if we reset the value of an intermediate output following the max-pooling layer, it is as if we deactivated the contribution of that specific filter in the final result. This property is particularly important because it allows to evaluate the importance of each filter on the final prediction. Suppose we want to determine the contribution of code A123. First, we need to search the structure containing FILTER-CODE matches and see if the code A123 is present for one or more filters. Suppose that the chosen code has the following matches: A123-F1, A123-F2, A123-F3. At this point, the intermediate results (those after the max-pooling layer) corresponding to these three filters are set to zero. By performing the prediction without these events in the EHR, it is possible to determine how much the predicted values change, determining the contribution of code A123. Repeating this procedure for each code, we obtain the interpretability of the whole EHR.
The contribution of a single code is evaluated based on the following strategy: how much does the prediction result change by zeroing out the internal activations related to the considered code? Suppose that the EHR we choose has a prediction with Label 1 = 90%. If, to evaluate the contribution of the code A123, the prediction with Label 1 drops to 60%, it means that this code has a high contribution on this Label. To quantify the contribution of a code we can compute the difference of these two probabilities, in this case 30%. This does not mean in any way that such code has a contribution on the entirety of the prediction equal to 30%, but only that removing the contribution of the code internally to the model, the prediction of Label 1 diminishes by 30%. The reported values must be considered as an indicative value that allows us to understand the reasons of such prediction, preventing to use the model as a black box. Fig. 6, 7 and 8 visually describe the aforementioned concepts.
Figure 8: Implementation of the model-dependant interpretability procedure – third step.
In Fig. 9, we report the result of our interpretability procedure applied to a real-world example. The two arrows in the image represent two groups of temporally distinct codes (i.e., two consecutive admissions), thus allowing the model to use the previous information and “weighing them temporally” thanks to the memory of the ConvLSTM layer. For each procedure and diagnosis, we report their importance on both the prediction of readmission (Label 1) and the prediction of not readmission (Label 0) within 30 days. Clearly, the results of our interpretability procedure provide hints regarding which codes most influence the prediction relative to a single admission.
Figure 9: Interpretability of the model on a 30 days readmission prediction with model-dependant interpretability procedure.
5.1 Model Validation
@@
So far we focused on just one admission; however, to confirm the validity of the predictions of our models, it is necessary to obtain the average results over many admissions so that we can understand which procedures and diagnoses have a greater contribution on average. The same interpretability framework can be used to validate the overall quality of the model considering all admissions. In this section, we present the interpretability results for both tasks of readmission, within 30 and 180 days. In Fig. 10 we show the results obtained. Average results are reported for Label 1 (readmission) for both tasks. For brevity, we report the 20 most common codes for each category.
@@
Analyzing the results of the 30-days readmission task, we see in first position the ICD-9-CM code related to the insertion of a central venous catheter related to several procedures in interventional and diagnostic/therapeutic settings (PROC 3897). Liver disease, including portal hypertension (DIAG 5723) and alcoholic cirrhosis (DIAG 5712), also appears to have an important weight in short-term readmission. Overall, the model has a high propensity for cardiovascular and respiratory diseases. We can see that the most represented cancers are breast cancer (DIAG V103)—first cancer in terms of prevalence in women—and malignant prostate cancer (DIAG V1046)—first in terms of prevalence in men.
Analyzing the results of the 180-days readmission task, we see among the most common manifestations cardiac (DIAG 42831, DIAG 42821), cerebral (DIAG 3484, DIAG 4321), hepatic (DIAG 5723, DIAG 5712), but also the injection of chemotherapeutic drugs for the treatment of neoplasms (PROC 9925). Overall, codes related to the cardiac/neurological domain are frequently present, and we can also note the importance attributed to multiple sclerosis (DIAG 340), which has a known fluctuating course with exacerbations.
In summary, comparing the 30 and 180 days readmission tasks for both models, we can see that the cardiological field is important both in the short and long term, as well as the hepatic field. Invasive diagnostic/interventional procedures, as well as pulmonary problems are present in 30-day readmission, while neurology is relevant only in 180-day readmission.

### v2

@@ 3.2 Readmission Prediction Model
Our model is based on a neural network that receives as input two distinct vectors, the first containing the succession of patient hospital admissions (blue dashed box in Fig. 2), the other containing the corresponding time gaps at each admission (green dashed box in Fig. 2). The core of our solution is represented by the Embedding layer and the ConvLSTM architecture.
@@
The Embedding layer receives as input a patient medical record for a specific admission in the form of temporally divided ICD-9-CM codes, as represented in Fig. 1. At this point, each code is converted into the corresponding unique word embedding vector, allowing us to better represent their semantics. This layer is necessary since computers work with numbers and cannot process strings in their native format. To improve the performance of the system, we opted for a transfer learning approach. Specifically, the pre-trained word embedding was obtained by [41], who used the Word2Vec model with SkipGram implementation [39] to train on healthcare records of 4 million patients over a time span of 2-4 year. The embedding vectors are 300-dimensional and contain many medical code standards, in our case we use only ICD-9-CM related to diagnoses and procedures. By means of dimensionality reduction techniques such as t-SNE [50], it is possible to represent the vectors generated by the word embedding model in 2 dimensions. This technique allows to graphically represent the semantically similar vectors as two neighboring points. Fig. 3 shows the result produced by t-SNE, with cosine similarity as evaluation metric to estimate the semantic relatedness. In Fig. 3 we show ICD-9-CM codes of both diagnoses and procedures, colored in blue and orange respectively. As evident from the graph, the interaction space of diagnoses and procedures is very complex, but it is possible to identify clusters defined by a specific medical domain.
@@
After each code has been converted into the corresponding embedding vectors, the ConvLSTM architecture implements the forecasting algorithm. This deep learning framework, introduced by [51], is an extension of LSTM, in which the matrix multiplication is replaced with the convolution operation. As a result of its structure, this innovative architecture allows to exploit both the spatial and temporal information of the input, characteristics necessary for our system to consider both long-term dependencies between the admissions and abstract general health information (e.g., comorbidity) from sequences of specific procedures and diagnoses. Following the ConvLSTM layer, max-pooling is performed, with the goal of making data depend only on the number of filters applied in the convolution and not on the initial input size. The second input contains a vector of categorical data (green dashed box in Fig. 2), equal in size to the time frames considered. The categories are those related to the various time gaps. Such data, being a simple sequence of mono-dimensional vectors, are processed by an LSTM layer, which permits to give the right weight to each event passed in the sequence. Subsequently, the outputs coming from the two branches of the network are concatenated (red dashed box in Fig. 2) and then a classifier with two Dense layers is used to predict whether a readmission will take place or not.
4 Interpretability Approach
Recent advancements in deep learning techniques have led to state-of-the-art results across various domains. However, one of the most significant challenges associated with deep learning is the interpretability of their outcomes [13, 12]. These models are often regarded as black boxes, lacking internal transparency due to their complex network structure, which consists of multiple layers performing non-linear transformation. In critical fields such as medicine, the interpretability of predictive tasks is of paramount importance, as a single decision can have profound implications on a patient’s life. A first attempt at improving the interpretability of deep learning models for readmission prediction was introduced by [15], based on the examination of convolution filters. These filters act as n-gram detectors, which can be interpreted as “clinical motifs”. As discussed earlier in this paper, identifying clinical motifs is essential for accurately predicting unplanned readmissions. However, as shown by [47], interpreting these filters can be challenging, as a single filter may capture multiple n-grams and potentially cancel the contributions of others. In this context, a popular framework applicable to machine learning and deep learning models is LIME [46], LIME approximates complex models with simpler, interpretable models in the local region of the prediction, allowing users to understand the factors influencing individual predictions. In the literature, methods such as LIME are considered among the most viable way to improve interpretability in neural networks [28].
For this reason, we initially applied LIME to determine the contribution of each code in the EHR to the overall prediction. This method works by perturbing the input data based on the concept of “local fidelity”, offering a straightforward approach to identify the key factors influencing a specific prediction. However, this method does not yield satisfactory interpretability results. The challenge lies in the specific structure of the ConvLSTM1d model input, which consists of two temporally structured vectors—–one for ICD-9-CM codes and the other for time gaps—that do not allow for easy perturbation. As a result, we developed a custom model-dependent interpretability approach designed to provide clearer insights when EHRs include temporal information. The goal of this approach is to highlight the importance of each code in the final prediction, thereby assisting medical professionals in understanding the factors that most influence the outcome. As highlighted by [48], this approach is feasible only because our prediction model’s architecture consists of few layers, thus greatly improving its interpretability. This would not be as effective with more complex architectures. The goal of this approach is to obtain an easy-to-understand output structured as follows:
  1. 1.
    Patient’s EHR, consisting of the various associated ICD-9-CM codes.
  2. 2.
    The prediction results, in conjunction with the probability of each label.
  3. 3.
We now describe the approach step by step. First, consider the output of a generic layer of the neural network: the correspondence between the output and the originating code is lost due to the max-pooling layer, which selects the largest response from each filter. To obtain an interpretable result, we need to keep track, for each filter, of the input code that originates the largest response. This allows to obtain a structure containing, for each input code, the correspondence between a filter and an ICD-9-CM code, named FILTER-CODE. It may happen that a code is not the maximum response of any filter, therefore not all the codes will be present in the FILTER-CODE structure. The second step is to modify the intermediate results coming from the max-pooling layer. The results of the max-pooling layer are necessarily $\geq 0$, since this layer is preceded by a ReLu activation function. Consequently, by resetting the value of an intermediate output following the max-pooling layer, we “deactivate” the contribution of that specific filter in the final result. This property is particularly important to evaluate the importance of each filter on the final prediction.
Figure 4: Implementation of the model-dependent interpretability approach – first step.
Figure 5: Implementation of the model-dependent interpretability approach – second step.
For instance, to determine the contribution of code A123, we first search the structure containing FILTER-CODE matches to identify whether code A123 appears in one or more filters. Suppose the following matches: A123-F1, A123-F2, A123-F3. At this point, the intermediate results (i.e., the resutls after the max-pooling layer) corresponding to these three filters are set to zero. By performing the prediction without these events in the EHR, it is possible to determine how much the predicted values change, thereby determining the contribution of code A123. By repeating this procedure for each code, we obtain the interpretability of the whole EHR.
To summarize, the contribution of a single code is evaluated with the following strategy: how much does the prediction change by zeroing out the internal activations related to the considered code? Suppose the EHR has a prediction with Label 1 = 90%. Intuitively, if the prediction with Label 1 drops to 60% when evaluating the contribution of the code A123, it means that A123 has a high contribution on this Label. To quantify the contribution of a code we compute the difference of the two probabilities, i.e., $90-60=30\%$. The reported contribution values must be considered as an indicative value that allows us to understand the reasons of such prediction, preventing a blind use of the model. Fig. 45 and 6 visually describe the approach step by step.
Figure 6: Implementation of the model-dependent interpretability approach – third step.
Fig. 7 shows the result of the interpretability approach applied to a real-world example. The two arrows represent two groups of temporally distinct codes (i.e., two consecutive admissions), thus allowing the model to leverage the previous information and “weigh them temporally” as a result of the memory of the ConvLSTM layer. For each procedure and diagnosis, we report their importance on both the prediction of readmission (Label 1) and the prediction of no readmission (Label 0) within 30 days. The results of this approach provide hints regarding which codes have the greatest influence on the prediction for a single admission. By computing the average contributions across multiple admissions, it is possible to understand which procedures and diagnoses contribute most to the overall predictions. Furthermore, through the use of charts and visualizations (e.g., Fig. 10), the model predictions can be easily explained to less technical individuals, such as clinicians.
Figure 7: Interpretability of the model on a 30 days readmission prediction with model-dependent interpretability approach.
@@ 5 Experiments and Results
In this section, we first describe the training data used for ConvLSTM1d, then we report and discuss the performances on two prediction tasks: readmission within 30 and 180 days. The main objective is to verify that our model improves performances in predicting unplanned hospital readmissions compared to state-of-the-art models, while at the same time allowing the interpretability of the results. Moreover, we compare ConvLSTMd1 with Deepr [15] (a CNN-based model for readmission prediction) and with two traditional machine learning models, namely logistic regression and random forest. The latter models are reported since they are widely used in hospital readmission prediction [52]. For logistic regression and random forest, we use a bag-of-words model instead of word embeddings. Bag of words is a simple yet popular approach in NLP problems, where text is represented as a set of words, without considering semantics or word order [53].
@@ 5.1 Training Dataset
The EHRs used to train ConvLSTM1d are taken from the MIMIC III Database, or simply MIMIC (Medical Information Mart for Intensive Care) [37]. MIMIC is a large, freely-available database comprising anonymized healthcare data associated with over 40,000 patients who stayed in intensive care units. The initial pre-processing of the dataset is fundamental to avoid errors in the data that may propagate to the subsequent steps. During pre-processing, we take inspiration from the procedure described by [10], that also use MIMIC III data for a task of readmission prediction.
@@
Fig. 8a and 8b report the data distributions from which we gathered this information. As a result, the two datasets we are going to use have the following characteristics:
@@
Additionally, the two generated datasets are balanced. Otherwise, as evidenced by Fig. 8c and 8d, the negative label would be predominant in both tasks (with a prevalence of 94% and 87% respectively), and accuracy could not be used as a metric. In both datasets, there may be multiple admissions for a single person. The presence of multiple admissions is important to extrapolate as much information as possible from the rows of the datasets with a positive label (minority label).
@@ 5.2 Training Parameters
After splitting the datasets into training, validation and test sets, we train our model using Adam optimizer and binary cross-entropy loss function. Table 1 reports the hyper-parameters used for the two tasks of 30-days and 180-days admission. Fig. 9 represents the learning curves for both loss and accuracy on the 30-day unplanned readmission task.
@@
5.3 Performance Metrics
Before presenting the performance obtained by ConvLSTMd1, we carry out a brief summary of the main metrics adopted. In the following description, TP stands for True Positive, TN for True Negative, FP for False Positive, FN for False Negative. Accuracy is an intuitive metric that consists in the ratio between the correctly predicted observations and the total number of observations. To obtain a good validity, the dataset must be balanced. Accuracy is an intuitive metric that measures the ratio of correctly predicted observations to the total number of observations. It is most informative when the dataset is balanced. Accuracy is useful to give an overall sense of performance.
$Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$
Precision is the ratio of correctly predicted positive observations to the total predicted positive observations. It is particularly useful when the cost of False Positives is high, since it focuses on the quality of positive predictions, minimizing the risk of incorrect positive classifications.
$Precision=\frac{TP}{TP+FP}$
Recall is the ratio of correctly predicted positive observations to the total number of actual positive observations. It is particularly useful when the cost of False Negatives is high, since it favors capturing as many positive observations as possible, even at the risk of including False Positives.
$Recall=\frac{TP}{TP+FN}$
F1 Score is the harmonic mean of precision and recall, providing a balanced measure that accounts for both False Positives and False Negatives. It allows the balance between precision and recall, offering a single metric that reflects both the quality and completeness of positive predictions.
$F_{1}=2\cdot\frac{\mathrm{p}\cdot\mathrm{r}}{\mathrm{p}+\mathrm{r}}$
where $p$ stands for Precision and $r$ for Recall.
@@ 5.4 Results and Discussion
Table 2 illustrates the performance results of ConvLSTMd1 for both 30-days and 180-days unplanned readmission prediction. Our model is trained to achieve the highest accuracy, which is used as the main evaluation metric. To give a better overview, we compare the results of ConvLSTMd1 with two traditional machine learning approaches for data prediction (logistic regression and random forest) and with Deepr [15], the most similar and advanced solution in literature for predicting hospital readmissions from EHRs. All results are obtained from carrying out 10 separate training sessions for each task and considering the average score between sessions. This expedient is necessary since deep learning models tend to obtain slightly different performance scores even for the same task.
@@
From Table 2 it is evident that ConvLSTM1d is the best model in terms of accuracy, substantially outperforming all the other models. In particular, logistic regression and random forest obtain good results in either precision or recall for certain specific cases, but with overall accuracy and F1 scores significantly lower compared to our model. Deepr considerably improves on both accuracy and F1 scores with respect to traditional machine learning methods, but it prevails over ConvLSTM1d only in recall for unplanned readmission within 30 days. However, it should be noted that precision is much lower. Consequently, this is reflected in the F1 score, which favors ConvLSTM1d in both tasks. On the other hand, ConvLSTM1d obtains higher results in both accuracy and F1 scores for both labels and both tasks.
5.5 Model Interpretability Validation
@@
In this section, we apply the interpretability approach introduced in Section 4. To validate the predictions made by our models, it is crucial to understand which procedures and diagnoses contribute most to the predictions on average. To achieve this, we compute the average contributions across multiple admissions. A similar approach can be used to evaluate the overall quality of the model by averaging the results across all admissions. In this section, we present the results of the validation approach for both the readmission tasks considered in this paper, i.e., readmission within 30 and 180 days. Fig. 10 shows the results obtained. We report the average results for Label 1 (readmission) in both scenarios. For the sake of brevity, we present the 20 most common codes in each category.
Analyzing the results of the 30-days readmission task, in first position is ranked the ICD-9-CM code related to the insertion of a central venous catheter related to several procedures in interventional and diagnostic/therapeutic settings (PROC 3897). Liver disease, including portal hypertension (DIAG 5723) and alcoholic cirrhosis (DIAG 5712), also appears to have an important weight in short-term readmission. Overall, the model has a high propensity for cardiovascular and respiratory diseases. The most represented cancers are breast cancer (DIAG V103)—first type of cancer in terms of prevalence for women—and malignant prostate cancer (DIAG V1046)—first type of cancer for men. Considering the 180-days readmission task, among the most common manifestations are cardiac diseases (DIAG 42831, DIAG 42821), cerebral diseases (DIAG 3484, DIAG 4321), hepatic diseases (DIAG 5723, DIAG 5712), but also the injection of chemotherapeutic drugs for the treatment of neoplasms (PROC 9925). Overall, codes related to the cardiac and the neurological domain are frequently present, and it can also be noted the importance attributed to multiple sclerosis (DIAG 340), which has a known fluctuating course with exacerbation. In summary, by analyzing the results of both 30 and 180 days readmission tasks, we can conclude that the cardiological field is important both in the short and long term, as well as the hepatic domain. Invasive diagnostic and interventional procedures, as well as pulmonary problems, are very much present in the 30-day readmission scenario, while neurology is relevant only in 180-day readmissions.